OpenStudy (anonymous):

y varies inversely as x. When y = 5, x = 12. Determine y when x = 15.

OpenStudy (anonymous):

Is it 4?

OpenStudy (anonymous):

5*12=y*15 so y=4 you're right :)

thomaster (thomaster):

i think 12/5 = 2.4 so thats the factor 5 * 2.4 = 12 ? * 2.4 = 15 (15*5)/12 = 6.25 or \[6\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

given y=k/x,where k is a constant when x=12 ,y=5 5=k/12, or k=60 y=60/x when x=15,y=60/15=4